CAVITY RINGDOWN SPECTROSCOPY OF THE DARK Ã ^{2}E′′ STATE OF NO_{3}

\underline{A. DEEV}, J. SOMMAR and M. OKUMURA, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, California 91125.

The forbidden Ã ^{2}E′′ - X̃ ^{2}A′ transition of NO_{3} has been observed in the near IR (1.1-1.5 micron) by cavity ringdown spectroscopy. Several new vibronically allowed bands are recorded, and definitive vibrational assignments and rotational contour analysis will be presented, allowing for a more precise estimate of the band origin. The influence of the Jahn-Teller effect on the transition structure will be discussed.